A professional painter can paint 828 ft squared per hour. How do i convert this to yd squared per minute?

Answer by Theo(13342)

Painter can paint 828 square feet per hour.

to convert 1 hour to equivalent minutes, you need to multiply 1 hour by 60 to get:

the ratio of 828 square feet per hour becomes:

828 square feet per 60 minutes.

This is the same as the fraction of 828/60.

If you want to make the denominator equal to 1 minute, then you have to divide the numerator and the denominator by 60 to get:

13.8/1 which is the equivalent of:

13.8 square feet per minute.

828 square feet per 60 minutes is equivalent to 13.8 square feet per minute.

Now you want to convert square feet to square yards.

1 square yard equals 1 yard by 1 yard.

1 yard is equivalent to 3 feet.

1 square yard is therefore equal to 3 feet by 3 feet = 9 square feet.

To convert square feet to yards, you need to divide the square feet by 9 to get the equivalent square yards.

13.8 square feet is equivalent to 1.53333333 square yards.

Your final answer is that 828 square feet per hour is equivalent to 1.533333333 square yards per minute.

To confirm, work your way back.

1.533333333 square yards * 9 = 13.8 square feet per minute.

Multiply numerator and denominator by 60 to get:

13.88 square feet per minute is equivalent to 828 square feet per 60 minutes.

Since 60 minutes is equivalent to 1 hour, this means that:

828 square feet per 60 minutes is equivalent to 828 square feet per hour.

the major conversion factors used were:

1 hour = 60 minutes
1 yard = 3 feet
1 square yard = 1 * 1 yards
1 square yard = 3 * 3 feet = 9 square feet